OpenStudy (anonymous):

Solve −2x2 − 16x − 44 = 0. x equals negative 8 plus or minus I square root of 6 x equals negative 8 plus or minus 2i square root of 6 x equals negative 4 plus or minus i square root of 6 x equals negative 4 plus or minus 2i square root of 6

OpenStudy (anonymous):

Here

OpenStudy (anonymous):

I'm pretty sure it is c But help?

pooja195 (pooja195):

Thats correct

pooja195 (pooja195):

\[\huge~\rm~~x=\frac{ -b~\pm \sqrt{b^2-4ac} }{ 2a } \] \[\huge~\rm~~x=\frac{ -(-16)~\pm \sqrt{(-16)^2-4(-2)(-44)} }{ 2(-2) } \]

pooja195 (pooja195):

aww man it got cut >.<

pooja195 (pooja195):

http://prntscr.com/7k71uw

pooja195 (pooja195):

there c:

OpenStudy (anonymous):

\[x^2+8x+22=0\] \[x^2+8x+16-16+22=0,x^2+8x+16=-6\] \[\left( x+4 \right)^2=-6,x+4=\pm \sqrt{-6},x=-4\pm \sqrt{-6}\]